How much of a 10 liter 30% acid solution must be replaced with pure acid to obtain 10 liters
of a 50% solution?

Answer by TimothyLamb(4379)

---
30(10 - x) + 100x = 50*10
300 - 30x + 100x = 500
300 + 70x = 500
70x = 200
x = 2.8571429
---
answer:
replace 2.857 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php